40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15996     Entry time: Mon Apr 5 22:26:01 2021
Author: gautam 
Type: Update 
Category: LSC 
Subject: PRMI 1f locking (no ETMs) recovered 

Since it seems like the entire electronics chain has no obvious failure, I decided to compensate for the apparent increased optical gain by turning the flat whitening gain down from +18dB to 0dB. Then, after some fiddling around with alignment, settings etc, I was able to lock the PRMI once again, with the ETMs misaligned, using REFL55_I to sense PRCL, and REFL55_Q to sense MICH. Some sensing matrices attached. Some notes:

  1. I made some changes to the presentation so that the units of the sensing matrix are now in [W/m] sensed on a photodiode. 
    • The info printed on the plot is also more verbose, I now indicate explicitly the actuator driven to make the measurement, and also the drive frequency. The various gains used to convert counts to Watts on a photodiode are also indicated.
    • I thought about printing the actual matrix too but haven't arrived at a clean prez style yet.
    • This is to facilitate easier comparison to Finesse models / analytic calcs.
    • I take into account all the gains from the photodetector to the servo error point where the measurement is made. However, the splitting fractions between various photodiodes is not included, so you will have to do that yourself when comparing to a Finesse model.
    • For example, in pg2 of Attachment #1, the measured gain of PRCL sensed in REFL55_I is ~2e6 W/m. But only ~4% of IFO REFL ends up on the REFL55 photodiode. So, this measurement would be consistent with a Finesse simulated optical gain of ~50MW/m, which is in the ballpark of what I do actually see.
  2. There seems to be reasonable agreement between Finesse and these measurements. But why should the old settings / locking have worked at all then?
  3. I tried two schemes for MICH actuation today.
    • The first was the usual BS + PRM combo, and I got the sensing matrix on pg 1 of Attachment #1. With this scheme, the MICH/PRCL orthogonality is a joke.
    • Then I changed the MICH actuator to the ITMs, and got the sensing matrix on pg 2 of Attachment #1. With this scheme, the orthogonality looks much better. I think the slight non-orthogonality in the 11/33 MHz photodiodes may even be reasonable, since the 11 MHz field isn't a good sensor of the anti-symmetric modes, but have to confirm by calculation/simulation. But certainly the separation of signals is much cleaner when the ITMs are used to control MICH.

So there is clearly something funky with the nominal MICH actuation scheme (MICH suspension, PRM suspension or both), which we should get to the bottom of before trying any low noise locking. I think using the ITMs as the MICH actuator in the full lock will not be a good low nosie strategy, as we would then be "polluting" all our suspended optics with our control loops, which seems highly suboptimal for technical noise sources like coil driver noise etc.

Attachment 1: PRMI_Apr5sensMat_consolidated.pdf  363 kB  Uploaded Mon Apr 5 23:38:21 2021  | Hide | Hide all
PRMI_Apr5sensMat_consolidated.pdf PRMI_Apr5sensMat_consolidated.pdf
ELOG V3.1.3-